OpenStudy (anonymous):

Prove that: tanx(cot x+tan x)=sec^2x

OpenStudy (anonymous):

Using the fact that : \[\cot(x) = \frac{1}{\tan(x)}\] when we multiply that out we get: \[\tan(x)(\frac{1}{\tan(x)})+\tan^{2}(x) = 1+\tan^{2}(x) = \sec^{2}x\]

OpenStudy (anonymous):

tanx(cot x+tan x) = 1 + tan^2 x = sec ^2x

OpenStudy (anonymous):

the\[1+\tan^{2}(x) = \sec^{2}(x) \] comes from this equation: \[\sin^{2}(x)+\cos^{2}(x) = 1\] and dividing everything by cos^2(x)